Using the formula for the length of the arc, compute s, r or θ which ever is missing.

a. s = 6π r = 8

b. s = 10π r = 8


1
Expert's answer
2021-01-26T04:18:54-0500
"s=r\theta""\theta=\dfrac{s}{r}"

a. "s=6\pi, r=8"


"\theta=\dfrac{6\pi}{8}=\dfrac{3\pi}{4}"

b. "s=10\pi, r=8"


"\theta=\dfrac{10\pi}{8}=\dfrac{5\pi}{4}"